### v2

# RadioActive: 3D Radiological Interactive Segmentation Benchmark

Abstract
Current interactive segmentation approaches, inspired by the success of META’s Segment Anything model, have achieved notable advancements, however, they come with substantial limitations that hinder their practical application in 3D radiological scenarios. These include unrealistic human interaction requirements, such as slice-by-slice operations for 2D models on 3D data, a lack of iterative interactive refinement, and insufficient evaluation experiments. These shortcomings prevent accurate assessment of model performance and lead to inconsistent outcomes across studies. The RadioActive benchmark overcomes these challenges by offering a comprehensive and reproducible evaluation of interactive segmentation methods in realistic, clinically relevant scenarios. It includes diverse datasets, target structures, and interactive segmentation methods, and provides a flexible, extendable codebase that allows seamless integration of new models and prompting strategies. We also introduce advanced prompting techniques to enable 2D models on 3D data by reducing the needed number of interaction steps, enabling a fair comparison. We show that surprisingly the performance of slice-wise prompted approaches can match native 3D methods, despite the domain gap. Our findings challenge the current literature and highlight that models not specifically trained on medical data can outperform the current specialized medical methods. By open-sourcing RadioActive, we invite the research community to integrate their models and prompting techniques, ensuring continuous and transparent evaluation of interactive segmentation models in 3D medical imaging.
1 Introduction
Figure 1: Current interactive segmentation methods require clinicians to interact with radiological images slice-by-slice, leading to increased workload.
Accurate segmentation of anatomical structures or pathological areas is crucial in fields like radiology, oncology, and surgery to isolate affected regions, monitor disease progression, and treatment planning, and guide therapeutic procedures. On the one hand, traditional supervised medical segmentation models have demonstrated strong performance across a range of anatomies and pathologies [18, 19, 17, 37]. However, their applicability remains heavily constrained by the limited task-specific training data available, as well as their deterministic predictive behavior that cannot be influenced, possibly requiring labor-intensive manual post-processing of predictions. Consequently, fully autonomous AI solutions have not yet found widespread autonomous clinical applications.
On the other hand, numerous semi-automatic segmentation techniques, not reliant on AI, are already in clinical practice to expedite manual annotation processes [13]. These current ad hoc methods do not yet tap into the potential of AI-based automation to drastically reduce annotation time. A method that bridges this gap and allows clinicians to segment any target with just a single click within the image could greatly enhance the efficiency of clinical workflows.
Interactive segmentation methods, like the Segment Anything Model (SAM) [20] represent exactly this. SAM is designed to segment any target through various user interaction methods, including point-based and bounding box prompts. This allows users to easily specify the area of interest by clicking on it or drawing a bounding box around it, making the segmentation process both flexible and intuitive. A particularly powerful feature is the ability for users to iteratively refine initial predictions by adding more positive or negative prompts.
Due to their great potential, interactive segmentation methods have attracted a lot of attention in the medical domain and led to many studies evaluating and adapting SAM for 3D medical image segmentation [34, 5, 15, 45, 30, 3, 24, 11]. Moreover, several researchers have been inspired by SAM’s capabilities to develop their own methods, often specifically designed for the 3D nature of radiological data [9, 12, 21, 39].
Although these domain-specific adaptations to medical data have shown promising progress, many published methods are plagued by pitfalls that obfuscate the efficacy of the models and prevent clinicians and researchers from determining the best methods for their use cases:
Figure 2: Some models operate natively in 3D and enable full 3D interaction. Only models that accept mask prompts allow iterative refinement of initial predictions with human guidance.
Applying interactive 2D models to 3D data on a slice-by-slice basis (P1): Assuming clinicians will interact with each slice individually is unrealistic and undermines the efficiency improvements these methods aim for. Moreover, a slice-by-slice approach introduces an unfair bias when comparing 2D and 3D models, as 3D models typically require only a few interactions per image, leading to significantly fewer interactions and less supervision [3, 24, 43, 42, 41, 27].
Neglecting refinement (P2): Many studies assess interactive segmentation methods based on a single interaction step, overlooking the inherent ambiguities in radiological images [24, 9, 11, 2]. Often, a second interaction may be necessary to specify which specific substructure the clinician wants to segment. This could be, e.g. a vessel within the liver, or the necrosis within a tumor, as exemplified in the well-known BraTs segmentation challenge [4]. Furthermore, clinicians often want to adapt the segmentations to their clinic’s local protocol or refine them, particularly for targets with high inter-rater variability, like pathological structures [10, 1, 14]. Overall, there is a notable lack of research exploring realistic, iterative interaction methods for 2D models applied to 3D volumes.
Obfuscated and insufficient evaluation (P3): Currently, there is a lack of a standardized approach to evaluation. Often results are not incomparable between papers, with evaluation often obfuscated or insufficient. We observed the following shortcomings: (i) Not specifying whether predictions were interactively refined or based on a single prompt with multiple points [3, 39]. (ii) Being intransparent on the number of initial prompts given [9]. (iii) Using the best mask rather than the final mask after interactive refinement [39]. (iv) Evaluating predictions slice-by-slice or on sub-patches of a 3D volume instead of evaluating on the full image [34, 24, 3, 12, 21]. (v) Excluding targets considered too small, hence neglecting valid targets such as small lesions that are neither tested nor trained on [24, 3, 39]. (vi) Comparing against non-promptable models and SAM, rather than any other promptable model trained on medical data [3, 24, 11, 12]. (vii) Lastly, overemphasizing segmenting healthy structures, such as organs, where existing supervised public models already perform well [40, 37], instead of focusing on pathologies, where interactive refinement could provide the greatest benefits [39, 43].
To overcome these challenges, establishing a benchmark is essential, aligning with the recent recommendations by Marinov et al. [26]. To this end, we present RadioActive, a robust, reproducible, and extendable Radiological InterActive Benchmark for 3D medical image segmentation. This benchmark highlights the most effective 2D and 3D interactive segmentation approaches and tests realistic prompting methods in the 3D domain. We conduct evaluations directly in the 3D domain and across multiple models, providing more generalizing insights, as opposed to prior work [16, 26, 34]. Our key contributions include:
  1. 1.
    RadioActive, for the first time, enables a fair comparison of the most influential 2D and 3D interactive segmentation methods for 3D medical data. By measuring the number of simulated interactions, a proxy for the ’Human Effort’, we test different prompting strategies that do not require slice-by-slice interactions (P1).
  2. 2.
    We propose effective interaction strategies for refinement of predictions in a 3D volume, without requiring clinicians to interact with each individual slice (P2).
  3. 3.
    We provide a standardized evaluation protocol to generate prompts, select model outputs and compute the segmentation metrics on the entire image across ten datasets, covering various modalities and target structures, including small lesions (P3). Our benchmarking efforts include a performance comparison against leading interactive segmentation methods in the medical domain.
  4. 4.
    The extendable RadioActive framework allows developers to a) easily evaluate a new method in a fair manner against established methods and b) easily develop and investigate new prompting strategies.
Through open-sourcing RadioActive, we invite researchers to integrate their methods into our framework, promoting continuous and equitable assessment that allows us to track the overall progress in the field of interactive 3D medical image segmentation reproducibly and transparently.
Figure 3: RadioActive overview. Although our evaluation is performed on entire 3D volumes, the benchmark accommodates both 3D and 2D interactive segmentation methods. While 3D model prompting is relatively straightforward, we introduce prompting and refinement strategies for 2D models that minimize the effort required from human interaction. The benchmark is designed to be extensible, and researchers are encouraged to propose and integrate additional methods seamlessly using our codebase particularly for areas marked by three dots.
2 RadioActive
The RadioActive Benchmark is designed to easily enable a fair and reproducible evaluation of 2D and 3D interactive segmentation methods for 3D radiological image segmentation for the very first time Fig. 3. While prompting 3D models is generally straightforward, we introduce specific prompting and refinement strategies for 2D models to streamline human interaction and reduce the simulated Human Effort. The proposed benchmark includes seven established models and ten datasets covering different target structures and image modalities. All datasets are publicly available and we provide an automatic download and pre-processing for improved usability and reproducibility.
Moreover, the benchmark is built with flexibility in mind, enabling seamless integration of additional methods. Researchers are invited to contribute new approaches, particularly new models, new prompting schemes, and new interesting datasets to the collection. Overall, the design of our benchmark allows for easy testing and validation of novel segmentation methods, making the benchmark a catalyst for advancing methodology for interactive 3D medical image segmentation. In the following, we present the different components of RadioActive.
2.1 Initial Prompting
Prompts are a key component of any interactive segmentation method and can highly influence the achieved performance of the underlying method. RadioActive distinguishes between two visual prompting types. Point prompts correspond to a click of a user in the image, and box prompts refer to a box around the target structure. Notably, some methods also enable a distinction between foreground and background point prompts. While 3D models allow segmenting a 3D volume natively, 2D-based models require an interaction for each slice, resulting in excessive effort, which is prohibitive for clinicians as it would take too much time in daily clinical practice. Hence, any meaningful benchmark must account for this difference in prompting effort. To increase the feasibility of 2D models for 3D applications, it is essential to reduce this effort. We propose two straightforward methods, for both point and box prompts, to explore their performance and provide a proxy for measuring the effort of human interaction.
Point interpolation: Let $I\subseteq N$ be the set of axial indices of all foreground slices. We simulate a user by selecting $n$ foreground points, specifically the center of the largest connected component of slice $i_{1},...,i_{n}\in I$ where the $i_{j}$ are equally spaced within $I$ and $i_{1}=\min(I)$ and $i_{n}=\max(I)$. Then, we interpolate linearly between each point and the next one and use the intersections of the resulting lines with the axial slices as positive point prompts, as visualized in Fig. 4 c).
Point propagation: We simulate a user providing $\min(I)$, $\max(I)$, and a 2D point prompt within the slice corresponding to the median axial index $i_{m}$. Given this point, the model generates a segmentation $S_{m}$ for the median slice. We then calculate a central point, specifically the center of mass of the largest connected component of $S_{m}$, to use as a point prompt for the slice indexed by $i_{m-1}$. Again, we generate a segmentation $S_{m-1}$ of this slice, and create a new central point until we segment the slice with the axial index $\min(I)$. The propagation is then repeated upwards, starting from $i_{m+1}$ and continuing until we segment the slice with the axial index $\max(I)$. This process is visualized in Fig. 4 e).
Box interpolation: We simulate a user providing $n$ 2D bounding boxes, one in each of $i_{1},....,i_{n}\in I$, with $i_{j}$ defined as in the point interpolation paragraph. Since the boxes are uniquely defined by their minimum and maximum vertices, we can interpolate between the minimum vertices as in point propagation to get a minimum vertex in each axial slice, and similarly get a maximum vertex in each axial slice, providing a box prompt in each slice. This box interpolation is exemplified in Fig. 4 d).
Box propagation: We simulate a user providing $\min(I)$, $\max(I)$, and a 2D box prompt within the slice corresponding to the axial index $i_{m}$, with $m$ defined as in point propagation. The model then generates a segmentation $S_{m}$ for the median slice. A new bounding box is created based on $S_{m}$ and used as a prompt for the slice indexed by $i_{m}-1$. The propagation is continued down to $\min(I)$ and then repeated upwards until $\max(I)$ as in point propagation, but using box prompts instead of point prompts. See Fig. 4 f) for a visualization.
While realistic prompting behavior is the priority RadioActive also supports the previously used unrealistic slice-by-slice prompting styles for completeness. Additional details on prompt schemes are provided in the appendix in Sec. 6.
Figure 4: Different promoting schemes for 2D models based on point prompts (on the left) and box prompts (on the right). While a) and b) expect unrealistic human slice-by-slice interaction, c) and d) illustrate the proposed prompt interpolation schemes, where a human needs to provide prompts for at least 3 slices (4 slices in this case). Prompts for the remaining slices are generated by interpolating between the initial prompts. e) and f) present the proposed prompt propagation methods, where the prompt for each subsequent slice is automatically generated based on the model’s prediction from the previous slice. Only the initial slice and upper and lower boundaries require manual prompts.
2.2 Refinement Prompting
Refinement of previous segmentations is an important aspect of interactive segmentation models, as it allows iteratively improving the segmentation until the desired structure is segmented to a user’s demands. Some interactive segmentation models allow for the refinement of initial segmentations by providing the model with the previous prediction along with a new prompt to correct errors, either through foreground clicks on false negative pixels or background clicks on false positive pixels. While this process is straightforward for 3D models, 2D models naively only allow for refinement in a slice-by-slice fashion, which again places an unrealistic burden on clinicians. Therefore, we present refinement strategies that require a reasonable level of Human Effort.
Scribble refinement:
To represent a user-centric refinement strategy we introduce an algorithm simulating user-created scribble prompts: At each refinement step, our proposed algorithm generates either positive or negative additional prompts. The decision to generate positive prompts follows a Bernoulli trial with success probability $p=n_{fn}/(n_{fn}+n_{fp})$, where $n_{fn}$, $n_{fp}$ represent the number of false negatives and false positive voxels, respectively.
If positive prompts are selected, we perform a connected component analysis on the false negative voxels. Given $L$, the largest connected component, we generate a scribble from the bottom to the top of $L$ by taking the centroid of $L$ in each slice to simulate drawing a vertical scribble through the middle of $L$. This simulates a clinician annotating regions that were erroneously not segmented.
For 2D models, we then individually feed all slices $i\in I$ where the voxel along the scribble was not predicted, along with the new positive prompt derived from the scribble and the previous prediction $s\subset S$, back into the model. For 3D models, we feed the whole 3D patch, together with the previous prediction $S$, and all new positive points derived from the scribble into the network in one step.
If negative prompts are selected, we identify a non-axial slice $S_{fp}$ of $S$ that contains the most false positives. Then we generate a contour curve around the ground truth target object at a distance of 2 pixels. We then select a subpart $C$ with a length of 60% of the full curve and sample all pixels $c\in C$ that are false positives to obtain a set of points $D$, simulating a user drawing a few scribbles in areas where the model over-segmented the target. For 2D models, we then generate new slice predictions for each slice containing a point in $D$ by providing the model with the previous prediction as well as new negative prompts: all $d\in D$ which belong to that slice. For 3D models, we again feed the whole 3D patch, together with the previous prediction $S$ and a negative prompt sampled from $D$. The motivation for the different approaches for FPs is, that correcting the model near the target boundary is likely to provide more informative feedback. As none of the evaluated methods features a dedicated scribble prompt encoding, we instead generate a set of point prompts from our scribble refinement trajectory, which is similar to the recently proposed work of Wong et al. [41] for scribble prompting of SAM. Aside from this refinement strategy, we also support unrealistic refinement prompts for which details are provided in the appendix in Sec. 6.
2.3 Human Effort Proxy
A model’s performance is highly dependent on the effort a human invests into the initial prompting and the refinement of the predicted masks. Generally, the effort of using 3D methods is lower than the effort of using 2D methods, although the strategies mentioned above significantly reduce the effort of 2D methods.
We aimed to establish a quantitative measure of the effort a method would require from a human user, however, various issues arise: Considering a formalized mathematical approach where effort is quantified through the degrees of freedom (DoF) of each interaction. A point would correspond to 3 DoF, a 2D box 5 DoF (requiring selection of the z-axis and two 2D points), and a 3D box corresponds to 6 DoF. The point interpolation corresponds to 9 DoF, whereas point propagation would only be 5 DoF, since it requires just the axial coordinate rather than both minimum and maximum points with 3 DoF each. From the user’s perspective, however, identifying the z-coordinate demands the same level of effort as selecting a 3D coordinate by clicking at the target structure’s endpoint along the z-axis. Similarly, an arbitrary scribble has significantly more DoF than a straight or parabolic line, yet the difference in effort for the user is minimal. Subsequently, this approach was discarded.
Instead, we define user effort in terms of the number of interactions (i.e. clicks of a user) required for a specific prompt. This leads to e.g. a 2D and 3D click being equivalent in cost. While this is not an exact measure, it offers the most pragmatic estimation of the actual effort involved from a user’s perspective.
2.4 Interactive Segmentation methods
In our benchmark, we include various interactive segmentation methods. Fig. 2 illustrates the types of prompts each method supports. Iterative refinement is only possible for methods that allow a (previously predicted) mask as a prompt.
SAM is the most prominent model from the natural image domain, that inspired many researchers to evaluate and adapt it to the domain of radiological medical images. It was trained on iteratively generated and curated 1B masks and 11M images, but not explicitly on radiological images. It was the first to popularize interactive segmentation models [20].
SAM2 is an extension of SAM that was trained on even more images and introduced support for video data [33].
MedSAM is an adaptation of SAM that fine-tuned SAM’s weights on 1,570,263 image-mask pairs from the medical domain. It supports only a single forward pass without refinement and is limited to box prompts [24].
SAM-Med 2D is another adaptation of SAM, fine-tuned on 4.6 million images with 19.7 million masks from the medical domain. Unlike MedSAM, it supports points, boxes, and mask prompts, allowing for refinement [3].
SAM-Med 3D incorporates a transformer-based 3D image encoder, 3D prompt encoder, and 3D mask decoder. It was trained from scratch using 22,000 3D images and 143,000 corresponding 3D masks and supports point and mask prompts and also allows for refinement [39].
SAM-Med 3D Turbo is an updated version of SAM-Med 3D trained on a larger dataset collection of 44 datasets for improved performance. It supports the same prompt styles as SAM-Med 3D [39].
SegVol is an interactive 3D segmentation model based on a 3D adaptation of a ViT [7] that was trained on 96K unlabelled CT images and fine-tuned with 6K labeled CT images. It supports points and bounding boxes as spatial prompts but does not allow iterative refinement [9].
Aside from these models there exist other notable interactive models, such as Vista3D [12], 3D Sam Adapter [11] and Prism [21]. However, while being promptable, they are closed-set, i.e. not trained to segment any arbitrary prompted structure. Subsequently, they were not considered for this benchmark.
2.5 Datasets
Dataset selection was a non-trivial problem for this benchmark: While models that were originally introduced in the natural image domain rarely see any radiological 3D data, the medical counterparts were often trained on all publicly available datasets that the authors could obtain. For example, MedSAM was trained using more than 60 publicly available datasets [24]. Although these methods conducted their final validation on excluded datasets or at least on separate test subsets of images, the test datasets vary between models. As a result, identifying annotated datasets with interesting target structures that were not part of any of the included methods’ training datasets has proven challenging.
Nevertheless, we assembled a diverse collection of ten lesser-known or recently released public datasets featuring various pathologies and organs, including CT and MRI image modalities. Specific details of these are provided in Tab. 1. While the quantity of cases may seem low, many contain multiple instances or target structures. Moreover, due to the pathological nature of many they represent the currently common size of datasets in the medical domain. To enhance reproducibility and eliminate barriers of entry for non-domain experts, we automated the dataset download and pre-processing, minimizing any required domain knowledge to use the benchmark. However, due to the sparsity of labeled datasets, we urge developers to exclude these datasets from their train dataset selection, as inclusion would compromise the integrity of a clean evaluation through RadioActive.
Table 1: Overview over all Datasets. None of these datasets were part of the original training data for the methods, except for SegVol, which utilized D2 HanSeg.
Dataset Modality Targets Images
D1 MS Lesion [31] MRI (T2 Flair) MS Lesions 60
D2 HanSeg [32] MR (T1) 30 Organs at risk 42
D3 HNTSRMFG [38] MRI (T2) Oropharyngeal cancer & metastatic lymph nodes 135
D4 RiderLung [44] CT Lung lesions 58
D5 LNQ [6] CT Mediastinal lymph nodes 513
D6 LiverMets [36] CT Liver metastases 171
D7 Adrenal ACC [29] CT Adrenal tumors 53
D8 HCC Tace [28] CT Liver, Liver tumors 65
D9 Pengwin [22] CT Bone fragments 100
D10 Segrap [23] CT 45 Organs at risk 30
2.6 Evaluation
All interactive segmentation methods identify their target structure based on a spatial prompt, inherently resulting in instance segmentation. As a result, we evaluate on an instance-by-instance basis. Unlike in object detection, each prompt already provides information on the localization of the target structure, making detection metrics like F1-Score irrelevant. Subsequently, we rely solely on the Dice Similarity Coefficient (DSC) score as a metric. The instance-wise DSC metric is then averaged per case (i.e. per image volume), and further aggregated across all cases in the dataset, as recommended by Maier-Hein et al. [25]. For better presentation, we averaged the DSC across all classes of a dataset and also specified how many human interactions were simulated.
3 Experiments and Results
In our experiments, we evaluate all seven models across all datasets. First, we evaluate the performance during the initial prompting step, and following this, we evaluate the iterative refinement performance of methods that support this. For each of these, we test models in both, realistic and high-effort prompting schemes. To enhance readability, we present the results together with the primary findings in the main manuscript, while additional, more granular insights are deferred to Sec. 7 in the appendix.
3.1 Initial prediction
Unrealistic effort:
As an upper baseline, we begin with an idealized and unrealistic scenario where each slice is prompted individually for all 2D models. In this setting, we evaluate different numbers of point prompts per slice (PPS), as well as alternating positive and negative prompts ($\pm$ PPS), and slice-wise box prompts with varying numbers of boxes per slice (BPS). Figure 5 shows that models employing box prompts achieved significantly higher average Dice scores, with SAM2 demonstrating the strongest performance across all models. Conversely, point-based prompts performed poorly, particularly for small target regions, such as small MS lesions in dataset D1 (see Table 2 in the appendix). Although including positive and negative prompts and increasing the number of point prompts led to improvements, these were minor compared to the marked superiority of box-based prompts. Interestingly, generalist models not trained with a particular focus on medical data performed worse with point prompts per slice but outperformed models trained on medical data when using box prompts. This may be because models trained with medical data are better at resolving the ambiguity of point prompts, especially given the often low contrast between pathologies and surrounding tissue. In contrast, SAM and SAM2 models excel in object identification when using box prompts.
Figure 5: Unrealistic prompting of 2D Boxes each slice performs best. 2D models prompted with one Box Prompt Per Slice (BPS), indicated with a star, outperform models prompted with various Point Prompts per Slice (PPS, line plots). Drawing a 2D box accounts for 2 interactions similar to providing 2 points. While alternating positive and negative points (dashed lines) is slightly superior to only positive points the gap between points and boxes remains large.
Realistic Effort:
To simulate a human-in-the-loop scenario, we evaluate various prompting strategies that avoid slice-by-slice interaction. As described in Sec. 2, for 2D models, we test point and box interpolation, as well as propagation, using different numbers of initial prompts. For 3D models, we explore varying numbers of Point prompts Per Volume (PPV) and 3D box prompts. Fig. 6 presents the following key findings:
  1. 1.
    For all models, simple prompt interpolation methods achieve similar performance as the corresponding prompt per-slice methods. For example for SAM-Med 2D, 3 Point Interpolation even outperforms prompting 1 PPS having a mean DSC of 26.08 compared to a DSC of 23.13. This is because we select central points for the interpolation methods instead of random foreground points. Another example is SAM prompted using 5 Box Interpolation, that has a mean DSC of 56.9 compared to a DSC of 57.84 when prompted with a box for each slice.
  2. 2.
    SAM 2 outperforms specialized medical models across all prompting schemes using box interpolation.
  3. 3.
    Among 3D models, only SegVol is competitive to 2D models that use box prompts.
  4. 4.
    Both box and point propagation perform worse than their interpolation counterpart, though this may improve as models evolve.
Figure 6: Simple Interpolation Strategies Match Unrealistic Slice-Wise Prompting. Sampling prompts from the interpolated connection between three initial prompts yields similar performance for SAM2 as slice-wise prompting (left). This is also observed for box interpolation across all models (middle). 3D models perform worse than 2D methods when only a few points are provided, while SegVol demonstrates that using a 3D box is superior to points (right).
3.2 Interactive Refinement
Finally, we evaluate the performance of the models during iterative refinement. For 2D models, this involves predicting on a slice-by-slice basis. As illustrated in Fig. 7 (top), adding refinement prompts to each slice results in a substantial performance boost. While the proposed scribble-based refinement consistently improves outcomes, it does not achieve the same level of improvement as adding a prompt to every slice, which is expected since not all imperfect slices are guaranteed to receive new prompts during the scribble refinements. Moreover, we observed that for 2D models, it is crucial to provide the initial prompts again for each of the refinement steps. In the appendix in Table 5, we present results without including the previous point in the iterative refinement step, which resulted in a performance decline during refinement. 2D models tend to over-segment the target, filling the entire slice foreground. The absence of the initial prompt leads to a complete loss of target location information, as the initial predicted mask is highly inaccurate. Our refinement likely generates negative additional prompts due to the large number of false positive pixels. Qualitative examples of this failure case are presented in the appendix in Sec. 7 in the appendix.
For 3D models, iterative refinement also led to consistent performance improvements. Both randomly sampled prompts and those derived from refinement scribbles improved performance with each refinement iteration. Although SegVol initially performs best for the initial prediction and achieves a mean DSC of 43.32 prompted with 5 centered point prompts, it lacks support for further refinement. In contrast, SAM-Med 3D Turbo - worse in the initial prediction with a Dice of 40.89 - surpasses SegVol after several refinement steps reaching a DSC of 47.4.
Figure 7: All models demonstrate significant improvements from iterative refinement. Dashed lines represent the use of the proposed scribble refinement. While the unrealistic scenario of one refinement point per slice yields better performance, the proposed scribble refinement consistently enhances results across iterations for 2D models.
4 Discussion and Conclusion
In this paper we introduced RadioActive  and with it, compared the performance of 2D and 3D interactive segmentation models in 3D medical imaging. We provide a holistic and transparent overview of the current state-of-the-art and highlight key findings that offer practical insights:
  1. 1.
    Realistic 2D prompting can match unrealistic prompting: Our introduced realistic prompting styles can match unrealistic 2D prompting methods, see Fig. 6. This unlocks 2D methods for actual 3D medical image workflows without any major performance penalties.
  2. 2.
    SAM 2 outperforms all medical 2D & 3D models: When prompting SAM2 through a realistic prompting scheme such as the Box Interpolation Sam 2 is superior to all medical 2D & 3D models. It even surpasses models after unrealistic iterative refinements.
  3. 3.
    Iterative Refinement is essential: The ability to iteratively refine segmentations significantly enhances model performance, particularly in challenging cases. Models that allow multiple rounds of corrections show better accuracy, making this feature crucial for clinical applications. For example, SegVol reached the highest performance in a static setting for point prompts, however, SAM-Med 3D Turbo is able to exceed SegVol given a few interactions, highlighting the importance of refinement.
  4. 4.
    Points fail for difficult and small structures. Contrary to claims in previous literature [3, 34], point-based methods fail, likely due to previous work training and evaluating their methods on simpler target structures [24, 3, 34].
  5. 5.
    Bounding boxes outperform points. Consistent with previous studies [34, 24, 9, 27], we find that Bounding boxes consistently outperform point-based inputs by providing better spatial context, which leads to improved segmentation accuracy, especially for complex structures in radiological images. Point-based prompts lack this context, resulting in poorer performance.
Limitations
While the RadioActive benchmark addresses many issues, our work has limitations. Currently, we only approximate Human Effort. While the pragmatic approach provides valuable insights into model performance in relation to effort, it falls short of capturing the full complexity and practical challenges of real clinical applications. As future work, a comprehensive study involving clinicians is essential to assess the efficacy of different prompting strategies in real-world environments. Such a study should not only evaluate segmentation performance but also measure the time required for annotation, offering critical insights into the practical feasibility and efficiency of these models in clinical practice. Currently, the benchmark includes ten datasets, encompassing a variety of pathologies and selected anatomical structures in MRI and CT images. We acknowledge that the quantity of images in these datasets is relatively small—a common limitation for most rare lesion datasets—all large, however, most larger publicly available datasets have already been utilized in training the methods making this challenging. Another limitation of this work is that, given the rapid development of new interactive segmentation methods, very recent approaches, such as Wong et al. [41], are not yet included. However, we consider this a strength, as it highlights the need for a fair and open benchmark to keep up with ongoing advancements [26]. The published codebase is designed to seamlessly integrate new prompting strategies and models, enabling rapid adaptation and extension.
Conclusion
The open-source nature of RadioActive  serves as a catalyst for interactive segmentation research in 3D medical imaging by enabling a fair and reproducible comparison between existing methods. Future work can build upon and extend this framework to investigate crucial questions in interactive segmentation. Through this, we believe RadioActive will contribute to improving real-world clinical application of interactive segmentation methods, facilitating a reduction of labor for medical professionals and accelerating meaningful clinical research.
Supplementary Material
5 Model specifications
5.1 SAM
SAM is compatible with multiple image encoders, particularly the ViT family from [8]. We used the default and best-performing model with ViT-Huge. To ensure high-quality inputs for the model, we performed slice-wise inference by extracting slices from the inplane-plane axis. Each slice was normalized by first clipping values outside the 0.5th and 99.5th percentile of the volume’s intensity distribution and then scaling the values to [0, 255]. The image was repeated three times along the channel axis to produce an RGB-like image. Internally, SAM resizes these slices to 1024 pixels for the longest side with the shorter side being padded to 1024 pixels if needed to maintain square dimensions. Finally, the images are normalized using the model’s pre-stored mean and standard deviation as suggested by the original implementation. Inference was restricted to slices containing foreground. After prediction, the slices were reassembled into a volume, inverse transformed to the original coordinate system, and metrics were computed in the original image space.
5.2 SAM2
SAM2 supports multiple image encoders, specifically the Hiera family of Ryali et al. [35]. We used the best-performing model, Hiera-L. We clip the intensity values of the volumes based on the 0.5th and 99.5th percentiles, extract each slice along the through-plane, and make the images RGB-like just as with SAM. The images are then rescaled to $1024\times 1024$ pixels and again normalized using the mean and standard deviation provided together with the pretrained weights. Aggregation and inverse transformation are then performed similarly to SAM.
5.3 MedSAM
To apply the model slice-wise, we slice the input volume as with SAM, and then clip each slice based on their 0.5th and 99.5th percentile values. The images are then made RGB-like by repeating thrice along a new channel-dimension, rescaled to $1024\times 1024$ pixels and then normalised to [0, 1]. Aggregation and inverse transformation are performed similarly as with SAM.
5.4 SAM-Med2D
To apply the model slice-by-slice, we slice the input volume as with SAM, and then clip each slice based on their 0.5th and 99.5th percentile values same as with MedSAM. The slices are then made RGB-like and converted to a [0, 255] scale as in SAM’s preprocessing. The slices are then standardized using a mean and standard deviation provided along with the model and resized to $256\times 256$ pixels. Aggregation and inverse transformation are performed similarly as with SAM.
5.5 SAM-Med3D
The model is 3D so no slicing is needed. The volume is respaced to $1.5\times 1.5\times 1.5$ mm and then clipped based on its 0.5th and 99.5th percentiles. SAM-Med 3D performs inference on a 128x128x128 crop. The crop is centered around our point prompt if there is only one point prompt passed, and around the centroid of our prompts if multiple points are passed simultaneously. For subsequent refinement steps, the crop remains unchanged. The predicted crop is inserted back in its correct position within the wider coordinate system and then respaced back to the original spacing so that evaluation takes place in the corresponding native image space.
5.6 SAM-Med3D Turbo
SAM-Med3D Turbo is an updated checkpoint for SAM-Med-3D and so we perform the same pre- and postprocessing.
5.7 SegVol
Intensity values are clipped by its 0.5th and 99.5th percentiles. The mean and standard deviation of the foreground voxels are used for zscore normalization. The values are then rescaled to a [0,1]. Finally, the volume is cropped to its foreground. A first ’zoom-out’ inference is performed on this image, followed by a ’zoom-in’ sliding window inference. The predicted volume is then transformed back to the original space and compared with the unprocessed ground truth to calculate metrics.
6 Supported prompting schemes
2D initial prompts:
For the 2D models, every foreground-containing slice needs an initial prompt. RadioActive supports the following 2D prompting schemes:
  • N PPS: For each foreground-containing slice, N random foreground points are selected.
  • N±$\pm$±PPS: For each foreground-containing slice, $\lceil N/2\rceil$ random foreground and $\lfloor N/2\rfloor$ background points are selected.
  • Box PS: For each foreground-containing slice, we select a bounding box around the foreground.
  • NP Inter: N points are used for the proposed point interpolation prompting scheme as explained in Sec. 2.1. At least 3 points are needed.
  • P Prop: The center point of the slice with index $I_{m}$ is selected, and passed to the model along with $min(I)$ and $max(I)$ with notation and method as explained in Sec. 2.1.
  • NB Inter: N 2D boxes are used for the proposed box interpolation prompting scheme as explained in Sec. 2.1. At least 3 boxes are needed.
  • B Prop: The 2D box from the slice with index $I_{m}$ is selected, and passed to the model along with $min(I)$ and $max(I)$ with notation and method as explained inSec. 2.1.
3D initial prompts
For 3D models, the entire volume is predicted in a single step, making the prompting process more accessible for human interaction.
  • N PPV: N random foreground points are selected from the volume.
  • N center PPV: We generate a set of slice points using the point interpolation method as in Sec. 2.1 and then select a subset o $N$ points spaced equally along the axial direction.
  • 3D Box: A simple 3D bounding box around the foreground is used.
2D iterative refinement
The refinement experiments always consist of an initial prompt, followed by iterative refinement prompts that are generated as explained in the following. For all refinement steps, a point prompt from a previous step can be included.
  • 1PPS Refine: For each slice, a refinement point prompt is selected randomly from the misclassified pixels, and passed either a false positive or false negative as appropriate from the false positive or false negative predicted volume.
  • Scribble Refine: This prompting scheme follows the proposed scribble refinement scheme for 2D models in Sec. 2.2. Only slices that receive a new refinement prompt are predicted again.
3D iterative refinement
Again, a point prompt from a previous step can be reused in all subsequent refinement steps.
  • 1 PPV Refine: For each refinement step, one positive or negative point prompt per volume is selected randomly from the misclassified pixels, and passed either a false positive or false negative as appropriate.
  • Scribble Refine: We generate a refinement scribble as we would for 2D iterative refinement, but then randomly select just one refinement point from the scribble.
All 2D and 3D bounding boxes were generated with perfect alignment around the target object. Since the impact of non-perfect bounding boxes has been extensively studied in previous works, we chose not to explore this additional degree of freedom [34, 27]. However, the open-source nature of our framework facilitates straightforward extensions to include a wider range of prompting scenarios.
7 Additional Results
* Refer to caption
Figure 8: Qualitative examples of failure-cases: A) We highlight issues of SAM-Med3D, which is limited to predict within a small field-of-view. Segmentations are truncated at the edges, due to the organ exceeding the boundaries of the prediction volume. B) 2D segmentation models sometimes default to predicting the entire abdomen, due to the ambiguity of what was queried.
Difference between MRI und CT data
Table 8 presents the DSC for prompting schemes simulating realistic Human Effort without refinement, averaged over MRI datasets (D1-3) and CT datasets (D4-10). While direct comparisons between modalities are not possible due to differing tasks and datasets, notable trends emerge. Some models exhibit consistent performance across modalities, such as SAM2, which shows minimal variation. In contrast, all 3D models display significant performance differences between MRI and CT. This disparity likely reflects the training data distribution; for instance, SegVol was exclusively trained on CT data, leading to subpar performance on MRI datasets. However, on CT-specific tasks, SegVol matches the performance of SAM and SAM2, despite its overall lower average performance across all datasets.
Previous prompts are essential for 2D model refinement
For 2D models, it is important to reintroduce the initial prompts at each refinement step. In Tab. 5, we demonstrate that omitting the initial prompt during iterative refinement results in a performance decline. This behavior arises because 2D models often over-segment the target, filling the entire slice foreground. This is visualized in Fig. 8 B. Without the initial prompt, these models lose crucial target location information, as the initial predicted mask is typically highly inaccurate. Consequently, the refinement process likely generates erroneous additional prompts due to the abundance of false-positive pixels, further degrading performance.
Previous prompts hinder 3D model refinement
For 3D models, Tab. 6 demonstrates that incorporating previous prompts does not enhance refinement performance beyond the first step. This observation aligns with the findings in Fig. 6 (right), where SAM-Med 3D and SAM-Med 3D Turbo show no improvement when provided with more than three prompts simultaneously.
Performance comparison between organs and pathological structures
Table 7 presents a performance comparison between pathological and organ targets. While a direct comparison is challenging due to variations in tasks and datasets, certain trends can be observed. Notably, 2D models exhibit greater difficulty in segmenting pathologies when prompted with points alone, likely due to the typically higher contrast of organs compared to pathological structures.
Challenges of 3D patch-based prediction for larger structures
Table 7 shows the DSC for liver segmentation (D8a), a common task included in the training datasets of all medical models. However, SAM-Med 3D and SAM-Med 3D Turbo encounter difficulties in predicting the entire liver within a single patch. In contrast, SegVol’s zoom-out/zoom-in inference approach effectively addresses this limitation, achieving a superior DSC of 91%. Such a failure-case is visualized in Fig. 8 A.
Prompter Model Interactions D1 D2 D3 D4 D5 D6 D7 D8 D9 D10 Average
1PPS SAM 1X 0.81 2.66 3.13 2.06 1.42 1.0 7.86 12.47 3.42 3.33 3.82
1PPS SAM2 1X 1.25 2.74 4.08 3.9 3.07 1.51 9.6 16.35 3.97 3.83 5.03
1PPS SamMed 2D 1X 10.72 26.3 24.55 28.83 22.76 9.71 30.83 31.24 26.05 20.31 23.13
2$\pm$±PPS SAM 2X 0.95 2.79 3.41 3.29 2.12 1.06 8.88 14.31 3.59 3.85 4.43
2$\pm$±PPS SAM2 2X 3.39 3.75 7.01 4.89 4.03 1.93 13.17 17.03 4.2 4.97 6.44
2$\pm$±PPS SamMed 2D 2X 11.88 28.09 27.62 33.29 24.07 11.25 32.54 31.5 28.27 21.27 24.98
2PPS SAM 2X 0.88 2.66 3.15 2.56 1.88 1.03 8.1 14.08 3.35 3.7 4.14
2PPS SAM2 2X 1.6 2.78 4.44 3.9 3.12 1.54 9.82 16.74 3.97 3.93 5.18
2PPS SamMed 2D 2X 11.63 29.9 28.9 32.88 24.38 11.34 37.84 36.41 32.51 22.39 26.82
3$\pm$±PPS SAM 3X 1.02 2.89 3.6 3.61 2.65 1.18 9.66 15.07 3.63 4.08 4.74
3$\pm$±PPS SAM2 3X 4.22 4.22 7.78 5.57 4.56 2.08 14.6 17.94 4.36 5.13 7.05
3$\pm$±PPS SamMed 2D 3X 12.83 31.8 31.72 36.4 25.85 12.62 40.42 36.91 34.12 23.29 28.6
3PPS SAM 3X 0.96 2.79 3.18 2.91 2.34 1.18 8.54 14.7 3.44 3.88 4.39
3PPS SAM2 3X 2.29 3.37 5.64 4.41 3.74 1.79 10.87 17.2 4.09 4.19 5.76
3PPS SamMed 2D 3X 11.95 31.34 31.8 33.61 24.9 12.15 43.26 40.16 35.62 23.07 28.79
5$\pm$±PPS SAM 5X 1.01 2.7 3.98 4.31 2.28 1.31 11.76 15.97 3.87 4.15 5.13
5$\pm$±PPS SAM2 5X 7.37 6.15 10.41 6.79 6.57 3.01 23.18 20.41 5.56 6.81 9.63
5$\pm$±PPS SamMed 2D 5X 13.95 34.18 35.63 39.39 27.19 14.01 46.49 41.05 37.83 24.5 31.42
5PPS SAM 5X 1.14 3.37 4.0 3.94 2.96 1.35 10.54 15.5 3.78 4.54 5.11
5PPS SAM2 5X 4.04 4.67 7.91 5.74 4.5 2.01 15.05 17.77 4.78 4.89 7.14
5PPS SamMed 2D 5X 12.26 32.17 34.55 33.08 25.24 12.66 50.88 45.0 38.34 23.22 30.74
10$\pm$±PPS SAM 10X 2.03 4.77 7.13 8.59 6.32 1.68 18.79 21.29 4.11 9.9 8.46
10$\pm$±PPS SAM2 10X 15.61 11.26 15.78 11.5 12.12 7.65 33.5 23.09 7.98 10.01 14.85
10$\pm$±PPS SamMed 2D 10X 15.02 35.62 39.67 42.49 28.27 15.95 52.98 46.17 40.97 25.26 34.24
10PPS SAM 10X 1.3 4.0 5.29 5.3 3.46 1.4 13.45 16.73 4.3 6.66 6.19
10PPS SAM2 10X 4.61 5.15 7.58 6.12 4.8 2.2 19.8 18.38 5.68 4.79 7.91
10PPS SamMed 2D 10X 12.03 31.1 36.39 30.38 24.22 13.24 58.07 51.41 38.98 22.1 31.79
Box PS SAM 2X 13.27 58.73 68.26 63.2 66.22 74.72 70.15 69.46 32.34 62.07 57.84
Box PS SAM2 2X 70.25 64.18 73.21 73.06 72.07 76.51 73.39 67.9 47.83 65.95 68.44
Box PS SamMed 2D 2X 28.2 46.08 57.04 64.45 46.44 41.91 62.07 63.59 45.62 41.51 49.69
Box PS MedSam 2X 40.63 47.39 55.5 60.4 45.73 46.43 67.75 70.23 48.82 46.13 52.9
Table 2: Experimental results simulating unrealistic effort of a clinician. This involves prompting each slice of a 3D volume. ’PPS’ and ’BPS’ represent points per slice or box per slice, respectively. ’X’ implies that each interaction is replicated for every slice, multiplying the clinician’s effort across the entire volume.
Prompter Model Interactions D1 D2 D3 D4 D5 D6 D7 D8 D9 D10 Average
3P Inter SAM 3 0.84 3.6 3.49 2.22 1.67 1.03 8.28 12.67 3.48 4.56 4.18
3P Inter SAM2 3 1.38 3.67 4.75 4.18 3.41 1.63 10.32 17.01 4.17 5.18 5.57
3P Inter SamMed 2D 3 11.61 29.08 28.0 34.33 26.0 10.8 36.29 35.8 25.69 23.21 26.08
5P Inter SAM 5 0.84 3.6 3.5 2.17 1.7 1.02 8.3 12.61 3.64 4.5 4.19
5P Inter SAM2 5 1.38 3.68 4.73 4.1 3.37 1.63 10.44 16.98 4.4 5.13 5.58
5P Inter SamMed 2D 5 11.71 29.54 28.46 34.42 26.17 10.88 36.81 35.65 28.27 23.31 26.52
10P Inter SAM 10 0.84 3.61 3.48 2.21 1.71 1.03 8.26 12.64 3.66 4.49 4.19
10P Inter SAM2 10 1.38 3.67 4.76 4.14 3.43 1.62 10.39 17.01 4.5 5.1 5.6
10P Inter SamMed 2D 10 11.74 29.92 28.45 34.71 26.24 10.97 36.67 35.1 29.66 23.33 26.68
5P Prop SAM 7 1.09 3.49 3.6 1.78 1.7 0.99 7.84 13.7 3.39 4.32 4.19
5P Prop SAM2 7 3.77 3.75 4.59 3.47 3.13 1.6 8.96 17.52 3.78 4.79 5.54
5P Prop SamMed 2D 7 10.87 19.33 14.18 17.91 17.09 7.7 27.25 31.84 12.86 13.52 17.26
B Prop SAM 4 0.89 27.98 37.2 37.59 36.38 38.49 43.29 35.47 11.38 30.4 29.91
B Prop SAM2 4 3.82 33.77 43.31 40.82 41.08 40.88 46.54 37.39 19.37 31.68 33.87
B Prop SamMed 2D 4 2.82 23.78 30.17 36.12 25.15 22.56 37.98 45.6 19.1 21.0 26.43
B Prop MedSam 4 2.97 20.96 22.75 22.76 23.38 23.81 28.55 29.89 7.16 19.89 20.21
3B Inter SAM 6 13.13 54.04 63.34 58.64 64.26 71.35 65.52 61.78 25.58 59.23 53.69
3B Inter SAM2 6 69.76 57.59 67.71 69.08 69.06 73.11 68.35 63.24 37.21 63.26 63.84
3B Inter SamMed 2D 6 27.83 42.22 52.7 60.66 44.76 38.63 55.31 55.81 33.46 38.66 45.0
3B Inter MedSam 6 40.14 40.66 46.18 48.16 41.41 40.75 54.42 54.98 30.47 40.63 43.78
5B Inter SAM 10 13.25 56.3 66.81 62.58 65.76 73.63 69.5 68.25 31.33 61.54 56.9
5B Inter SAM2 10 70.13 61.16 71.53 72.57 71.32 75.52 72.72 67.25 46.01 65.42 67.36
5B Inter SamMed 2D 10 28.11 44.52 56.04 64.01 46.04 41.24 60.67 62.32 43.44 40.91 48.73
5B Inter MedSam 10 40.55 44.91 52.64 57.67 44.75 45.0 64.6 66.88 45.59 44.75 50.73
10B Inter SAM 20 13.27 57.71 67.81 63.17 66.12 74.35 70.02 69.21 32.24 61.98 57.59
10B Inter SAM2 20 70.25 63.0 72.65 73.05 71.94 76.12 73.19 67.62 47.48 65.83 68.11
10B Inter SamMed 2D 20 28.19 45.51 56.73 64.4 46.36 41.75 61.89 63.3 45.22 41.4 49.48
10B Inter MedSam 20 40.63 46.53 54.89 60.16 45.59 46.15 67.17 69.54 48.69 45.92 52.53
Table 3: Experimental results simulating a realistic clinician’s effort. ’PPS’ and ’PPV’ represent points per slice or volume, respectively. ’B Prop’ and ’P Prop’ denote the introduced box and point propagation schemes, while ’B Inter’ and ’P Inter’ refer to the introduced box and point interpolation methods.
Prompter Model Interactions D1 D2 D3 D4 D5 D6 D7 D8 D9 D10 Average
1PPV SamMed 3D 1 1.92 10.95 21.19 29.12 13.28 16.66 56.06 48.55 42.81 23.63 26.42
1PPV SamMed 3D Turbo 1 5.03 18.34 30.08 18.38 10.75 34.93 28.51 19.26 39.26 17.06 22.16
1PPV SegVol 1 8.84 21.28* 31.49 25.49 2.45 32.66 61.77 52.46 29.09 25.85 29.14
1 center PPV SamMed 3D 1 2.07 12.15 24.11 26.9 15.11 19.64 72.66 53.24 50.4 26.51 30.28
1 center PPV SamMed 3D Turbo 1 5.18 27.34 46.07 34.46 15.91 46.38 82.98 59.49 63.83 26.98 40.86
1 center PPV SegVol 1 9.96 24.91* 38.49 31.36 3.17 33.92 71.01 50.67 28.21 30.73 32.24
2 center PPV SamMed 3D 2 1.87 11.51 23.15 24.45 13.14 18.19 71.11 45.59 40.95 25.36 27.53
2 center PPV SamMed 3D Turbo 2 5.33 26.62 45.71 33.26 15.75 46.77 84.88 70.5 69.56 27.49 42.59
2 center PPV SegVol 2 11.2 31.31* 47.51 58.45 11.57 52.36 75.08 52.66 33.4 32.45 40.6
3 center PPV SamMed 3D 3 1.77 11.29 22.45 23.41 12.19 17.31 70.31 46.24 36.77 24.85 26.66
3 center PPV SamMed 3D Turbo 3 5.16 26.48 43.81 30.93 15.28 46.56 85.75 68.52 69.67 27.19 41.94
3 center PPV SegVol 3 11.52 31.1* 50.08 57.51 18.76 53.46 73.04 58.22 45.52 33.95 43.32
5 center PPV SamMed 3D 5 1.73 10.95 21.86 21.29 11.21 16.04 66.47 43.95 34.78 22.73 25.1
5 center PPV SamMed 3D Turbo 5 5.09 25.75 42.74 26.85 14.67 46.5 85.92 64.7 70.11 26.6 40.89
5 center PPV SegVol 5 11.7 31.17* 49.08 52.47 25.4 52.85 61.6 45.66 44.16 33.48 40.76
10 center PPV SamMed 3D 10 1.72 10.54 20.87 20.78 10.4 14.54 61.01 41.78 27.3 20.16 22.91
10 center PPV SamMed 3D Turbo 10 5.06 24.9 40.04 21.51 13.1 46.03 86.05 61.5 68.76 25.45 39.24
10 center PPV SegVol 10 11.69 30.22* 45.24 47.32 26.43 51.68 42.04 26.69 41.88 32.96 35.61
3D Box SegVol 3 0.55 37.17* 68.11 69.72 63.21 50.13 89.95 79.98 72.13 49.45 58.04
Table 4: Experimental results simulating a realistic clinician’s effort. ’center PPV’ stands for Point Per Volume, that was sampled from the center of the target object. Sampling from the center performs better than sampling a random point. SegVol trained on CT images from the D2 HanSeg dataset, while we tested on MRI images from the same patient cohort.
Iterations Prompter Model Interactions D1 D2 D3 D4 D5 D6 D7 D8 D9 D10 Average
0 1PPS + 1PPS Refine SAM 1X/1X 0.81 2.66 3.13 2.06 1.42 1.0 7.86 12.47 3.42 3.33 3.82
1 1PPS + 1PPS Refine SAM 1X/1X 1.02 2.94 4.46 3.22 1.97 1.13 8.73 13.81 3.5 3.56 4.43
2 1PPS + 1PPS Refine SAM 1X/1X 1.11 2.97 4.92 3.66 2.5 1.41 10.0 15.55 3.77 3.82 4.97
3 1PPS + 1PPS Refine SAM 1X/1X 1.18 2.37 5.24 4.41 3.03 1.55 11.06 16.97 4.02 4.0 5.38
4 1PPS + 1PPS Refine SAM 1X/1X 1.18 2.03 4.64 4.95 3.33 1.62 11.88 17.7 4.16 4.38 5.59
5 1PPS + 1PPS Refine SAM 1X/1X 1.11 2.2 3.3 5.56 3.37 1.43 12.57 18.25 4.09 4.32 5.62
0 1PPS + 1PPS Refine SAM2 1X/1X 1.25 2.74 4.08 3.9 3.07 1.51 9.6 16.35 3.97 3.83 5.03
1 1PPS + 1PPS Refine SAM2 1X/1X 2.96 4.99 8.69 6.2 5.64 3.32 17.94 24.31 5.49 5.78 8.53
2 1PPS + 1PPS Refine SAM2 1X/1X 5.25 6.83 11.68 8.03 7.64 5.04 24.56 29.52 7.39 7.83 11.38
3 1PPS + 1PPS Refine SAM2 1X/1X 7.53 8.8 15.02 10.01 9.52 6.87 29.11 33.53 9.06 9.59 13.9
4 1PPS + 1PPS Refine SAM2 1X/1X 9.81 10.51 18.31 11.99 11.09 8.24 33.22 48.49 10.45 11.54 17.37
5 1PPS + 1PPS Refine SAM2 1X/1X 12.14 12.15 21.33 13.52 12.56 9.21 37.04 51.38 11.72 13.5 19.45
0 1PPS + 1PPS Refine SamMed 2D 1X/1X 10.72 26.3 24.55 28.83 22.76 9.71 30.83 31.24 26.05 20.31 23.13
1 1PPS + 1PPS Refine SamMed 2D 1X/1X 12.44 29.97 28.54 31.35 24.3 11.78 39.25 37.54 31.41 21.45 26.8
2 1PPS + 1PPS Refine SamMed 2D 1X/1X 12.49 31.19 30.25 32.28 24.65 12.41 43.74 40.46 33.49 22.07 28.3
3 1PPS + 1PPS Refine SamMed 2D 1X/1X 12.59 31.93 31.49 32.85 24.98 12.88 46.69 42.32 34.71 22.6 29.31
4 1PPS + 1PPS Refine SamMed 2D 1X/1X 12.78 32.51 32.59 33.31 25.33 13.31 48.82 43.68 35.59 23.05 30.1
5 1PPS + 1PPS Refine SamMed 2D 1X/1X 12.93 32.98 33.67 33.74 25.76 13.67 50.46 44.81 36.29 23.49 30.78
0 1PPS + 1PPS Refine * SAM 1X/1X 0.81 2.66 3.13 2.06 1.42 1.0 7.86 12.47 3.42 3.33 3.82
1 1PPS + 1PPS Refine * SAM 1X/1X 0.06 0.27 0.63 0.6 0.15 0.08 3.34 3.38 1.45 0.61 1.06
2 1PPS + 1PPS Refine * SAM 1X/1X 0.33 0.95 1.55 1.64 0.76 0.67 6.2 11.11 2.61 2.45 2.83
3 1PPS + 1PPS Refine * SAM 1X/1X 0.22 1.06 1.56 1.42 0.86 0.51 5.16 7.74 2.14 1.54 2.22
4 1PPS + 1PPS Refine * SAM 1X/1X 0.38 1.08 1.73 1.4 0.5 0.57 6.07 10.85 2.61 2.3 2.75
5 1PPS + 1PPS Refine * SAM 1X/1X 0.2 0.93 1.34 1.29 0.62 0.58 5.53 8.34 2.07 1.69 2.26
0 1PPS + 1PPS Refine * SAM2 1X/1X 1.25 2.74 4.08 3.9 3.07 1.51 9.6 16.35 3.97 3.83 5.03
1 1PPS + 1PPS Refine * SAM2 1X/1X 1.33 2.94 5.86 4.34 3.44 1.76 11.73 18.45 4.13 3.74 5.77
2 1PPS + 1PPS Refine * SAM2 1X/1X 1.17 3.02 6.24 4.62 3.53 1.64 11.58 17.89 4.3 3.66 5.77
3 1PPS + 1PPS Refine * SAM2 1X/1X 1.29 3.05 6.31 4.93 3.69 1.73 11.78 17.81 4.34 3.67 5.86
4 1PPS + 1PPS Refine * SAM2 1X/1X 1.31 3.22 6.57 4.97 3.79 1.75 11.82 17.5 4.37 3.84 5.91
5 1PPS + 1PPS Refine * SAM2 1X/1X 1.23 3.21 6.6 4.86 3.74 1.76 11.97 17.78 4.4 3.8 5.93
0 1PPS + 1PPS Refine * SamMed 2D 1X/1X 10.72 26.3 24.55 28.83 22.76 9.71 30.83 31.24 26.05 20.31 23.13
1 1PPS + 1PPS Refine * SamMed 2D 1X/1X 10.45 28.62 27.4 31.37 23.26 10.47 38.52 37.22 30.63 20.04 25.8
2 1PPS + 1PPS Refine * SamMed 2D 1X/1X 9.88 28.71 28.05 31.7 22.67 10.37 41.79 38.9 31.55 19.47 26.31
3 1PPS + 1PPS Refine * SamMed 2D 1X/1X 9.6 28.45 28.26 31.42 22.18 10.58 43.5 40.16 31.9 18.97 26.5
4 1PPS + 1PPS Refine * SamMed 2D 1X/1X 9.2 28.17 28.29 31.08 21.78 10.44 44.71 40.81 31.99 18.52 26.5
5 1PPS + 1PPS Refine * SamMed 2D 1X/1X 8.99 27.84 28.26 30.56 21.49 10.54 45.45 41.43 31.99 18.2 26.48
0 1PPS + Scribble Refine SAM 1/3 0.81 2.66 3.13 2.06 1.42 1.0 7.86 12.47 3.42 3.33 3.82
1 1PPS + Scribble Refine SAM 1/3 0.89 2.65 4.53 3.24 2.11 1.18 7.48 11.65 3.55 3.68 4.1
2 1PPS + Scribble Refine SAM 1/3 0.93 2.76 4.45 3.54 2.35 1.37 8.01 11.24 3.52 3.41 4.16
3 1PPS + Scribble Refine SAM 1/3 0.96 2.12 3.5 3.49 2.09 1.4 8.01 10.6 3.4 3.15 3.87
4 1PPS + Scribble Refine SAM 1/3 0.92 1.94 2.08 3.43 1.94 1.4 8.06 9.84 3.27 3.01 3.59
5 1PPS + Scribble Refine SAM 1/3 0.88 1.69 1.4 3.43 1.87 1.14 7.66 9.54 3.14 2.56 3.33
0 1PPS + Scribble Refine SAM2 1/3 1.25 2.74 4.08 3.9 3.07 1.51 9.6 16.35 3.97 3.83 5.03
1 1PPS + Scribble Refine SAM2 1/3 1.74 3.88 8.13 5.39 5.97 3.01 16.14 18.28 4.66 5.07 7.23
2 1PPS + Scribble Refine SAM2 1/3 2.18 4.42 10.05 5.84 7.57 4.14 19.28 19.15 4.86 5.75 8.32
3 1PPS + Scribble Refine SAM2 1/3 2.52 4.8 10.93 6.16 8.2 5.06 21.87 25.97 4.9 6.17 9.66
4 1PPS + Scribble Refine SAM2 1/3 2.83 5.03 12.37 6.26 8.75 5.88 24.46 27.84 4.86 6.45 10.47
5 1PPS + Scribble Refine SAM2 1/3 3.1 5.3 13.38 6.26 9.08 6.93 25.81 30.02 4.86 6.85 11.16
0 1PPS + Scribble Refine SamMed 2D 1/3 10.72 26.3 24.55 28.83 22.76 9.71 30.83 31.24 26.05 20.31 23.13
1 1PPS + Scribble Refine SamMed 2D 1/3 11.8 28.84 26.82 29.72 23.07 11.32 37.22 36.7 29.87 20.7 25.61
2 1PPS + Scribble Refine SamMed 2D 1/3 11.71 29.29 27.88 29.5 22.7 11.63 40.64 38.77 31.61 20.65 26.44
3 1PPS + Scribble Refine SamMed 2D 1/3 11.55 29.52 28.26 29.15 22.67 11.68 42.57 39.92 32.62 20.68 26.86
4 1PPS + Scribble Refine SamMed 2D 1/3 11.49 29.62 28.33 29.19 22.6 11.76 43.94 40.71 33.25 20.74 27.16
5 1PPS + Scribble Refine SamMed 2D 1/3 11.46 29.75 28.72 28.87 22.52 11.88 44.93 41.32 33.68 20.83 27.4
0 3B Inter + Scribble Refine* SAM 6/3 13.13 54.04 63.34 58.64 64.26 71.35 65.52 61.78 25.58 59.23 53.69
1 3B Inter + Scribble Refine* SAM 6/3 0.99 3.66 3.24 3.53 2.33 6.14 7.78 11.88 4.68 6.43 5.07
2 3B Inter + Scribble Refine* SAM 6/3 0.85 1.95 1.28 3.17 9.04 5.28 7.49 14.4 4.14 7.43 5.51
3 3B Inter + Scribble Refine* SAM 6/3 0.8 4.11 2.95 2.19 3.13 3.36 7.14 9.85 4.02 4.67 4.22
4 3B Inter + Scribble Refine* SAM 6/3 0.71 2.45 2.01 2.21 2.28 1.89 6.08 9.09 3.22 3.57 3.35
5 3B Inter + Scribble Refine* SAM 6/3 0.76 1.99 2.46 2.51 2.16 1.42 6.15 9.34 3.19 3.47 3.34
0 3B Inter + Scribble Refine* SAM2 6/3 69.78 57.59 67.71 69.08 69.06 73.11 68.35 63.24 37.21 63.27 63.84
1 3B Inter + Scribble Refine* SAM2 6/3 10.03 20.17 25.37 19.13 23.65 18.12 25.37 27.66 11.34 15.93 19.68
2 3B Inter + Scribble Refine* SAM2 6/3 5.07 6.98 10.53 7.31 8.35 3.99 20.7 19.58 5.61 6.41 9.45
3 3B Inter + Scribble Refine* SAM2 6/3 4.23 4.43 7.8 6.93 5.64 3.72 14.43 16.35 4.89 5.34 7.38
4 3B Inter + Scribble Refine* SAM2 6/3 3.99 3.84 6.96 7.22 4.86 3.53 13.37 15.02 4.56 4.9 6.83
5 3B Inter + Scribble Refine* SAM2 6/3 3.41 3.7 6.75 5.7 3.99 2.79 12.49 13.59 4.3 4.47 6.12
0 3B Inter + Scribble Refine* SamMed 2D 6/3 27.83 42.22 52.7 60.66 44.76 38.63 55.31 55.81 33.46 38.66 45.0
1 3B Inter + Scribble Refine* SamMed 2D 6/3 28.4 44.72 57.53 63.33 45.28 41.42 61.61 59.73 40.44 38.8 48.13
2 3B Inter + Scribble Refine* SamMed 2D 6/3 24.29 42.41 53.02 59.24 40.27 35.69 62.86 58.85 41.23 34.93 45.28
3 3B Inter + Scribble Refine* SamMed 2D 6/3 20.91 39.69 47.48 52.73 35.55 28.2 62.08 54.89 40.02 31.59 41.31
4 3B Inter + Scribble Refine* SamMed 2D 6/3 18.76 37.29 41.46 46.5 31.73 23.98 58.46 51.86 38.07 28.74 37.69
5 3B Inter + Scribble Refine* SamMed 2D 6/3 17.07 35.06 37.39 41.62 29.1 20.8 53.9 49.27 36.27 26.66 34.71
0 3P Inter + Scribble Refine SAM 5/3 0.84 3.6 3.5 2.17 1.7 1.02 8.3 12.61 3.64 4.5 4.19
1 3P Inter + Scribble Refine SAM 5/3 0.89 3.29 4.59 3.58 2.41 1.21 7.96 11.68 3.73 4.39 4.37
2 3P Inter + Scribble Refine SAM 5/3 0.98 3.19 5.18 4.21 2.65 1.44 8.64 11.54 3.68 4.24 4.58
3 3P Inter + Scribble Refine SAM 5/3 1.03 2.5 3.61 4.14 2.52 1.5 9.12 10.95 3.64 3.71 4.27
4 3P Inter + Scribble Refine SAM 5/3 1.03 2.04 2.23 4.14 2.31 1.48 8.91 10.4 3.48 3.32 3.93
5 3P Inter + Scribble Refine SAM 5/3 0.99 1.89 1.5 4.17 1.87 1.32 8.16 9.75 3.33 3.02 3.6
0 3P Inter + Scribble Refine SAM2 5/3 1.38 3.68 4.73 4.1 3.37 1.63 10.44 16.98 4.4 5.13 5.58
1 3P Inter + Scribble Refine SAM2 5/3 1.87 5.18 8.77 6.04 6.8 3.47 16.57 19.63 5.09 6.31 7.97
2 3P Inter + Scribble Refine SAM2 5/3 2.37 5.83 11.33 7.01 9.24 4.62 22.39 21.17 5.41 7.22 9.66
3 3P Inter + Scribble Refine SAM2 5/3 2.84 6.27 13.65 7.62 10.71 5.9 23.75 27.73 5.48 7.65 11.16
4 3P Inter + Scribble Refine SAM2 5/3 3.33 6.52 14.69 7.83 11.49 6.96 27.72 28.86 5.52 8.05 12.1
5 3P Inter + Scribble Refine SAM2 5/3 3.5 6.7 15.79 7.97 12.18 7.65 30.01 30.82 5.56 8.47 12.86
0 3P Inter + Scribble Refine SamMed 2D 5/3 11.71 29.54 28.46 34.42 26.17 10.88 36.81 35.65 28.27 23.31 26.52
1 3P Inter + Scribble Refine SamMed 2D 5/3 12.72 31.26 30.27 34.05 26.02 12.06 40.92 38.13 31.14 23.12 27.97
2 3P Inter + Scribble Refine SamMed 2D 5/3 12.46 31.66 30.58 33.13 25.57 12.19 42.79 39.25 32.55 22.88 28.31
3 3P Inter + Scribble Refine SamMed 2D 5/3 12.25 31.79 30.53 32.57 25.24 12.28 43.89 39.98 33.35 22.73 28.46
4 3P Inter + Scribble Refine SamMed 2D 5/3 12.13 31.83 30.42 32.12 25.05 12.33 44.89 40.52 33.84 22.63 28.58
5 3P Inter + Scribble Refine SamMed 2D 5/3 12.06 31.86 30.42 31.66 24.9 12.41 45.85 40.97 34.22 22.61 28.7
Table 5: Interactive refinement results for 2D models across 5 iterations. The initial prediction is made either using a single point per slice or one of our proposed prompting schemes. If the previous point prompt is not reused during refinement, indicated by a *, the performance drops. This also explains the drop in performance for the 3 Box interpolation after the initial prompt. The unrealistic slice-wise refinement (1 interaction per slice) is only slightly better than our proposed scribble refinement method (3 interactions).
Iterations Prompter Model Interactions D1 D2 D3 D4 D5 D6 D7 D8 D9 D10 Average
0 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 2.03 12.15 24.11 26.9 15.09 19.64 72.66 53.03 50.4 26.51 30.25
1 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.89 14.07 33.03 35.67 14.42 24.93 64.15 72.35 42.64 24.86 33.2
2 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.67 13.49 29.52 28.92 14.34 21.4 64.29 62.44 42.03 25.16 30.73
3 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.57 13.64 31.57 35.51 14.21 22.82 63.35 65.94 41.12 24.79 31.85
4 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.81 13.51 31.44 31.08 14.11 21.38 65.28 63.38 40.68 24.82 31.15
5 1 center PPV + 1 PPV Refine* SamMed 3D 1/1 5.46 13.56 32.19 33.49 14.18 24.28 62.12 65.2 40.41 25.12 31.6
0 1 center PPV + 1 PPV Refine* SamMed 3D Turbo 1/1 5.12 27.34 46.07 34.46 15.9 46.38 82.98 59.36 63.83 26.98 40.84
1 1 center PPV + 1 PPV Refine* SamMed 3D Turbo 1/1 2.81 21.83 27.6 17.69 8.25 30.88 79.91 72.82 61.07 21.37 34.42
2 1 center PPV + 1 PPV Refine* SamMed 3D Turbo 1/1 3.57 22.72 35.57 22.05 8.89 36.31 77.77 72.2 62.7 23.37 36.52
3 1 center PPV + 1 PPV Refine* SamMed 3D Turbo 1/1 3.25 22.22 32.87 17.88 8.74 33.27 77.73 66.93 61.51 22.17 34.66
4 1 center PPV + 1 PPV Refine* SamMed 3D Turbo 1/1 3.6 22.93 34.4 16.26 8.78 34.84 80.01 72.09 62.35 23.35 35.86
5 1 center PPV + 1 PPV Refine* SamMed 3D Turbo 1/1 3.09 21.72 32.32 14.83 7.66 31.79 76.55 70.69 62.1 22.19 34.3
0 1 center PPV + 1 PPV Refine SamMed 3D 1/1 2.03 12.15 24.11 26.9 15.09 19.64 72.66 53.04 50.4 26.51 30.25
1 1 center PPV + 1 PPV Refine SamMed 3D 1/1 3.16 13.08 27.85 36.99 14.07 23.65 72.64 78.47 49.24 26.72 34.59
2 1 center PPV + 1 PPV Refine SamMed 3D 1/1 4.12 12.99 28.79 36.47 12.5 24.22 72.88 82.11 48.78 26.71 34.96
3 1 center PPV + 1 PPV Refine SamMed 3D 1/1 4.56 12.97 28.68 35.59 11.36 24.71 72.23 82.5 49.52 26.54 34.87
4 1 center PPV + 1 PPV Refine SamMed 3D 1/1 4.9 13.11 27.59 36.86 10.75 24.55 72.48 82.87 49.31 27.01 34.94
5 1 center PPV + 1 PPV Refine SamMed 3D 1/1 4.94 13.04 27.36 36.85 10.35 25.03 73.23 82.92 49.42 26.95 35.01
0 1 center PPV + 1 PPV Refine SamMed 3D Turbo 1/1 5.12 27.34 46.07 34.46 15.9 46.38 82.98 59.37 63.83 26.98 40.84
1 1 center PPV + 1 PPV Refine SamMed 3D Turbo 1/1 5.65 28.81 48.08 38.1 16.41 50.17 86.5 73.12 67.78 28.59 44.32
2 1 center PPV + 1 PPV Refine SamMed 3D Turbo 1/1 5.82 29.45 48.88 43.04 16.94 51.87 87.34 79.33 69.56 29.79 46.2
3 1 center PPV + 1 PPV Refine SamMed 3D Turbo 1/1 5.86 30.04 49.11 46.83 17.56 53.0 87.71 80.79 70.36 30.81 47.21
4 1 center PPV + 1 PPV Refine SamMed 3D Turbo 1/1 5.97 30.56 49.35 45.66 18.3 54.12 87.78 82.91 71.25 31.6 47.75
5 1 center PPV + 1 PPV Refine SamMed 3D Turbo 1/1 6.17 30.92 50.24 49.17 19.24 54.74 88.15 84.43 71.94 32.13 48.71
0 1 center PPV + Scribble Refine SamMed 3D 1/3 2.03 12.15 24.11 26.9 15.09 19.64 72.66 53.04 50.4 26.51 30.25
1 1 center PPV + Scribble Refine SamMed 3D 1/3 3.31 12.58 25.65 32.52 13.57 23.05 71.32 68.61 47.72 26.04 32.44
2 1 center PPV + Scribble Refine SamMed 3D 1/3 4.03 12.93 25.8 34.42 11.66 24.54 72.38 78.9 47.89 26.2 33.88
3 1 center PPV + Scribble Refine SamMed 3D 1/3 4.42 12.93 26.36 34.38 10.89 24.88 73.54 84.12 48.88 25.95 34.64
4 1 center PPV + Scribble Refine SamMed 3D 1/3 4.37 13.05 26.78 35.61 10.29 25.33 73.91 85.34 49.14 25.81 34.96
5 1 center PPV + Scribble Refine SamMed 3D 1/3 4.56 13.23 27.32 36.3 9.85 25.76 73.7 86.4 49.55 25.87 35.25
0 1 center PPV + Scribble Refine SamMed 3D Turbo 1/3 5.12 27.34 46.07 34.46 15.9 46.38 82.98 59.37 63.83 26.98 40.84
1 1 center PPV + Scribble Refine SamMed 3D Turbo 1/3 5.36 27.71 47.84 37.92 15.93 48.96 86.18 71.33 67.78 27.93 43.69
2 1 center PPV + Scribble Refine SamMed 3D Turbo 1/3 4.84 28.19 48.07 40.22 16.6 50.74 87.2 76.94 69.93 29.01 45.17
3 1 center PPV + Scribble Refine SamMed 3D Turbo 1/3 4.34 28.75 47.42 42.91 17.08 52.24 87.71 78.53 70.86 30.08 45.99
4 1 center PPV + Scribble Refine SamMed 3D Turbo 1/3 4.22 29.41 47.72 43.91 17.65 53.4 88.06 80.09 71.9 31.06 46.74
5 1 center PPV + Scribble Refine SamMed 3D Turbo 1/3 4.3 29.84 48.23 45.84 17.89 54.62 88.24 80.88 72.54 31.66 47.4
Table 6: Interactive refinement results for 3D models over 5 iterations. The initial interaction always starts from a central point of the target object, and refinement is performed either by randomly sampling positive or negative points (1 interaction) or by selecting a point using the proposed scribble refinement method. Scribble drawing is counted as three interactions. In contrast to 2D models, including the previous point prompt did not improve the performance. It is indicated by a * if previous points are reused during refinement prompting.
Prompter Model Interactions D1 D2 D3 D4 D5 D6 D7 D10 D8a D8b Average Path. Average Org. Average
SAM NORM 3P Inter 3 0.84 3.6 3.49 2.22 1.67 1.03 8.28 4.56 18.98 6.37 5.1 3.41 9.05
SAM2 NORM 3P Inter 3 1.38 3.67 4.75 4.18 3.41 1.63 10.32 5.18 24.27 9.74 6.85 5.06 11.04
SamMed 2D NORM 3P Inter 3 11.61 29.08 28.0 34.33 26.0 10.8 36.29 23.21 40.98 30.62 27.09 25.38 31.09
SAM NORM 3B Inter 6 13.13 54.04 63.34 58.64 64.26 71.35 65.52 59.23 56.21 67.36 57.31 57.66 56.49
SAM2 NORM 3B Inter 6 69.76 57.59 67.71 69.08 69.06 73.11 68.35 63.26 57.43 69.05 66.44 69.45 59.43
SamMed 2D NORM 3B Inter 6 27.83 42.22 52.7 60.66 44.76 38.63 55.31 38.66 54.26 57.36 47.24 48.18 45.05
MedSam NORM 3B Inter 6 40.14 40.66 46.18 48.16 41.41 40.75 54.42 40.63 54.72 55.23 46.23 46.61 45.34
SamMed 3D NORM 1 center PPV 1 2.07 12.15 24.11 26.9 15.11 19.64 72.66 26.51 63.96 42.53 30.56 29.0 34.21
SamMed 3D Turbo NORM 1 center PPV 1 5.18 27.34 46.07 34.46 15.91 46.38 82.98 26.98 58.89 60.1 40.43 41.58 37.73
SegVol NORM 1 center PPV 1 9.96 24.91* 38.49 31.36 3.17 33.92 71.01 30.73 47.82 53.52 34.49 34.49 34.48
SegVol NORM 3D Box 3 0.55 37.17* 68.11 69.72 63.21 50.13 89.95 49.45 91.07 68.89 58.83 58.65 59.23
Table 7: Performance comparison for pathologies and organs. The table presents the average performance for all pathological datasets (D1, D3–D7, D8b) and all healthy organ datasets (D2, D8a, D19). Dataset D8 is split into two parts: D8a, which includes the liver, and D8b, which focuses on liver tumors as a pathological structure. Dataset D9, containing annotations of bone fractures, was excluded as it does not fit neatly into either category of healthy or pathological structures. * SegVol trained on CT images from the D2 HanSeg dataset, while we tested on MRI images from the same patient cohort.
Model Prompter Interactions Average MRI Average CT Average
MedSam 3B Inter 6 43.78 42.33 44.4
SAM 3B Inter 6 53.69 43.5 58.05
SAM2 3B Inter 6 63.84 65.02 63.33
SamMed 2D 3B Inter 6 45.0 40.91 46.76
SamMed 3D 3 center PPV 3 26.66 11.84 33.01
SamMed 3D Turbo 3 center PPV 3 41.94 25.15 49.13
SegVol 3 center PPV 3 43.32 30.9 48.64
SegVol 3D Box 3 58.04 35.28 67.8
Table 8: Averaged CT and MRI performance: The table shows the DSC for prompting schemes simulating a realistic Human Effort, without refinement, averaged over all MRI datasets (D1-3) and over all CT datasets (D4-10). The performance for models trained on medical data is highly dependent on the training data distribution: For example, SegVol was exclusively trained on CT data, leading to subpar performance on MRI datasets.